OpenStudy (bbb911):

Divide. Write the answer in lowest terms as a whole or mixed number if possible. 1/10 divided by 5/6 my answer was 3/5 and was just wanting to know if i got tht right

OpenStudy (anonymous):

1/10 // 5/6 = 6/50 = 3/25 Answer = 3/25

OpenStudy (bbb911):

what i did to get 3/6 was trying to get the common factor for 5 and 10 so i multiplied 6 by 2 and 5 by 2 i can't do that or ?

OpenStudy (bbb911):

i mean 3/5

OpenStudy (skullpatrol):

You need to invert and multiply.

OpenStudy (skullpatrol):

$\frac{ 1}{10} \div \frac{ 5}{6}$

OpenStudy (johnweldon1993):

exactly as @skullpatrol has said....1/10 divided by 5/6 take the reciprocal of the bottom fraction (flip it) then...just change the sign to multiplication and now you have 1/10 times 6/5 6/50 simplify farther

OpenStudy (skullpatrol):

$$\frac{ 1}{10} \div \frac{ 5}{6}$$

OpenStudy (bbb911):

yea i think my mistake was by trying to get the the common denominator to get 10 on the other fraction so i multiplied 6 and 5 by 2 so the fraction was like this 1/10 X 12/10 and went through simplifying i was just a bit confused

OpenStudy (skullpatrol):

You use "common denominators" for adding and subtracting fractions; not multiplying and dividing them.

OpenStudy (bbb911):

oooo okay tht help me a lot thank you two :)

OpenStudy (johnweldon1993):

anytime :)

OpenStudy (skullpatrol):

np :)